OpenStudy (anonymous):

what is the solution to the linear-quadratic system of equation y=x^2+5x-3 y-x=2

OpenStudy (zkoqmosii):

y = x^2 + 5x - 3 y - x = 2 set y - x = 2 equal to x, so x = 2-y. Then plug in x = 2 - y for the other equation such that: y = (2-y)^2 + 5(2-y) - 3 Then solve for y

OpenStudy (anonymous):

@zkoqmosii i got y=11-4y is that correct?